# literature.bib

### v2

@@ 2.2 Continuous time perspective
In order to analyze ULPDA, we want to define a similar time-continuous limit of (2.4). For that, we assume that the step sizes $\tau,\sigma$ are coupled via a parameter $\lim_{\tau,\sigma\to 0}\frac{\sigma}{\tau}=\lambda$. Taking both step sizes to zero at this ratio, the formal limiting Markov process satisfies the diffusion equation
$\left\{\begin{array}[]{l}\,\mathrm{d}X_{t}\in-(\partial g(X_{t})+K^{*}Y_{t})\,\mathrm{d}t+\sqrt{2}\,\mathrm{d}W_{t}\\ \,\mathrm{d}Y_{t}\in-\lambda(\partial f^{*}(Y_{t})-KX_{t})\,\mathrm{d}t,\end{array}\right.$ (2.5)
@@
$\,\mathrm{d}Z_{t}\in-A_{\text{cPD}}(Z_{t})\,\mathrm{d}t+B_{\text{cPD}}\,\mathrm{d}W_{t}.$ (2.6)
where we introduced the continuous-time primal-dual coefficients
$A_{\text{cPD}}(Z_{t})=\begin{pmatrix}\partial g&K^{T}\\ -\lambda K&\lambda\partial f^{\ast}\end{pmatrix},\qquad B_{\text{cPD}}=\begin{pmatrix}\sqrt{2}I\\ 0\end{pmatrix},$ (2.7)
with $I$ the identity in $\mathbb{R}^{d}$. Note in particular that the diffusion coefficient $\Sigma_{\text{cPD}}=\frac{1}{2}B_{\text{PD}}B_{\text{cPD}}^{T}$ is only positive semi-definite (or degenerate) with an $m$-dimensional null space, where $m$ is the dimension of the dual variable $Y$. As we will see in more detail, the corresponding mean-field limit PDE is not elliptic anymore, which complicates analysis of existence and uniqueness of solutions.
2.3 Spaces and notation
@@
$\mathcal{T}(\mu,\tilde{\mu}):=\inf_{\mathcal{M}\in\Gamma(\mu,\tilde{\mu})}\left(\mathbb{E}_{(x,\tilde{x})\sim\mathcal{M}}\left[\lVert x-\tilde{x}\rVert^{2}\right]\right)^{1/2}.$
where $\Gamma(\mu,\tilde{\mu})$ denotes the set of all couplings with marginals $\mu$ and $\tilde{\mu}$
@@
[…]
$\langle z,\tilde{z}\rangle_{a,b}=\langle x,\tilde{x}\rangle_{a}+\langle y,\tilde{y}\rangle_{b},\quad z=(x,y),\tilde{z}=(\tilde{x},\tilde{y})\in\mathcal{Z}.$
Accordingly, we define a weighted norm by
$\lVert z\rVert_{a,b}:=\sqrt{\langle z,z\rangle_{a,b}}=\sqrt{\lVert x\rVert_{a}^{2}+\lVert y\rVert_{b}^{2}},\quad z=(x,y)\in\mathcal{Z}.$
On the space $\mathcal{P}(\mathcal{Z})$, we will also denote $\mathcal{T}_{a,b}$ for the transport distance induced by the norm $\lVert\cdot\rVert_{a,b}$. Note that $\mathcal{T}_{a,b}$ is a metric on $\mathcal{P}_{2}(\mathcal{Z})$ and $\mathcal{T}_{1,1}=\mathcal{W}_{2}$. For every $a,b>0$, $\mathcal{T}_{a,b}$ is equivalent to $\mathcal{W}_{2}$, since the fact $\sqrt{\min\{a,b\}}\,d_{1,1}(z,\tilde{z})\leq d_{a,b}(z,\tilde{z})\leq\sqrt{\max\{a,b\}}\,d_{1,1}(z,\tilde{z})$ directly implies
$\sqrt{\min\{a,b\}}\,\mathcal{W}_{2}\leq\mathcal{T}_{a,b}\leq\sqrt{\max\{a,b\}}\,\mathcal{W}_{2}.$
3 The Stationary Solution in Continuous Time
We start by proving the existence of a unique stationary solution of the time-continuous limit (2.6) of the ULPDA algorithm. We further show that the corresponding primal-dual Fokker-Planck equation is contractive, implying convergence of any solution path to the stationary distribution. The contraction result is formulated for a general class of diffusion SDEs. It follows a standard calculation from the literature for overdamped Langevin diffusion and is generalized to the setting of an SDE with any strongly monotone operator as non-homogeneous drift coefficient.
@@ 3.1 Fokker-Planck Equation and Existence of a Stationary Solution
3.1.1 Contraction of Overdamped Langevin Diffusion via Coupling Argument
Suppose for now that the potential $h:\mathcal{X}\to\mathbb{R}$ is differentiable, that $\nabla h$ is $L$-Lipschitz-continuous and $\omega_{h}$-strongly convex. We repeat here an instructive calculation that has been carried out, e.g., in [Fournier2021, Durmus2019] to show stability of the overdamped Langevin diffusion process
@@ 3.1.1 Contraction of Overdamped Langevin Diffusion via Coupling Argument
In particular, if we choose one of the processes to be initialized at the target distribution as $\tilde{\mu}_{0}=\mu_{\ast}$, then the distribution of $\tilde{X}_{t}$ is stationary at $\tilde{\mu}_{t}=\mu^{\ast}$ and by the above we obtain
@@ 3.1.2 Contraction and Stationary Solution of the Primal-Dual Diffusion Equation
We can generalize the above argument allowing an analysis of the time-continuous limit (2.6) of the primal dual sampling scheme (2.4). For that, we consider $Z_{t}$ to satisfy a general stochastic differential inclusion with non-homogeneous coefficients
@@
[…]
@@
with $\Sigma=\frac{1}{2}BB^{T}\in\mathbb{R}^{n\times n}$.
@@
$\begin{pmatrix}\mathrm{d}Z_{t}\\ \mathrm{d}\tilde{Z}_{t}\end{pmatrix}\in-\begin{pmatrix}A(Z_{t})\\ A(\tilde{Z}_{t})\end{pmatrix}\mathrm{d}t+\begin{pmatrix}B\\ B\end{pmatrix}\mathrm{d}W_{t},$ (3.5)
@@
$\frac{\partial\Pi_{t}}{\partial t}=\nabla_{z}\cdot\left(w\pi_{t}\right)+\nabla_{\tilde{z}}\cdot\left(\tilde{w}\Pi_{t}\right)+(\nabla_{z}+\nabla_{\tilde{z}})\cdot(\Sigma(\nabla_{z}+\nabla_{\tilde{z}})\Pi_{t}),\qquad w\in A(z),\tilde{w}\in A(\tilde{z}).$ (3.6)
@@
By employing (3.6), we obtain with the same reasoning as before
@@
$=\frac{1}{2}\int\lVert z-\tilde{z}\rVert^{2}\left(\nabla_{z}\cdot\left(w\Pi_{t}\right)+\nabla_{\tilde{z}}\cdot\left(\tilde{w}\Pi_{t}\right)+(\nabla_{z}+\nabla_{\tilde{z}})\cdot(\Sigma(\nabla_{z}+\nabla_{\tilde{z}})\Pi_{t})\right)\,\mathrm{d}z\,\mathrm{d}\tilde{z}$
@@
Consider now again the primal-dual type diffusion (2.6) and denote $\pi_{t}$ for the law of $Z_{t}$. By Proposition 3.1, $\pi_{t}$ solves the Fokker-Planck equation
$\frac{\partial\pi_{t}}{\partial t}=\nabla_{z}\cdot\left(w\pi_{t}+\Sigma_{\text{{cPD}}}\nabla_{z}\pi_{t}\right),\qquad w\in A_{\text{{cPD}}}(z).$ (3.7)
with $\Sigma_{\text{{cPD}}}=\frac{1}{2}B_{\text{{cPD}}}B_{\text{{cPD}}}^{T}$ and $A_{\text{{cPD}}},B_{\text{{cPD}}}$ as in (2.7).
Since the drift coefficient $A_{\text{{cPD}}}$ is a monotone operator, the primal-dual type diffusion falls into the class of general diffusion equations considered in Lemma 3.3. Under strong convexity of $g,f^{\ast}$, the drift coefficient is strongly monotone and we obtain the following stability result.
@@
Assume that $g,f^{\ast}$ are $\omega_{g}$-strongly convex and $\omega_{f^{\ast}}$-strongly convex, respectively, with $\omega_{g},\omega_{f^{\ast}}>0$. Let $Z_{t},\tilde{Z}_{t}$ be two coupled processes that solve the primal-dual diffusion equation (2.5) with corresponding densities $\pi_{t}$ and $\tilde{\pi}_{t}$. Then it holds
@@
where $\omega=\min\{\omega_{g},\lambda\omega_{f^{\ast}}\}>0$. In particular, the dynamics are contractive and there exists a unique stationary distribution of the continuous time primal-dual diffusion process, denoted $\pi^{\text{{stat}}}_{\text{{cPD}}}$, which solves the primal-dual Fokker-Planck equation (3.7).
@@
We employ Lemma 3.3, hence the only part we need to show is the strong monotonicity of the negative drift coefficient $A_{\text{{cPD}}}$ defined by
$A_{\text{{cPD}}}=\begin{pmatrix}\partial g&K^{T}\\ -\lambda K&\lambda\partial f^{\ast}\end{pmatrix}.$
We write $z=(x,y)\in\mathcal{X}\times\mathcal{X}=\mathcal{Z}$ for the joint primal and dual variable and equip $\mathcal{Z}$ with the weighted inner product $\langle\cdot,\cdot\rangle_{1,\lambda^{-1}}$. Let now $w\in A_{\text{{cPD}}}(z)$ and $\tilde{w}\in A_{\text{{cPD}}}(\tilde{z})$. Then there are $u\in\partial g(x)$ and $v\in\partial f^{\ast}(y)$ such that $w=(u+K^{T}y,\lambda(v-Kx))$. Equivalently, if $\tilde{w}\in A(\tilde{z})$ then there exist $\tilde{u}\in\partial g(\tilde{x})$ and $\tilde{v}\in\partial f^{\ast}(\tilde{y})$ with $\tilde{w}=(\tilde{u}+K^{T}\tilde{y},\lambda(\tilde{v}-K\tilde{x}))$. It follows that
@@
The stationary solution of (3.7) will be denoted by $\pi^{\text{{stat}}}_{\text{{cPD}}}$ from here. We will also frequently refer to its $x$-marginal as $\mu^{\text{{stat}}}_{\text{{cPD}}}$.
@@ 3.2 Non-Concentration of the Stationary Solution
Having established the existence of $\pi^{\text{{stat}}}_{\text{{cPD}}}$, we can try to relate the algorithmic, time discretized scheme to it. However, while overdamped Langevin dynamics is designed to have the target $\mu^{\ast}$ as its stationary solution, this is not the case in the primal-dual setting. In general, $\mu^{\ast}$ is not the marginal distribution of $X$ when $(X,Y)\sim\pi^{\text{{stat}}}_{\text{{cPD}}}$, i.e. $\mu^{\ast}\neq\mu^{\text{{stat}}}_{\text{{cPD}}}$. We demonstrate this in a simple example of a Gaussian target, where we can solve the Fokker-Planck equation (3.7) in closed form.
@@
where $k\neq 0$ and $c_{g},c_{f}>0$. It holds $f^{\ast}(y)=\frac{c_{f}}{2}y^{2}$ for all $y\in\mathbb{R}$. The primal-dual flow (2.6) of the joint variable $Z_{t}=(X_{t},Y_{t})^{T}\in\mathbb{R}^{2}$ reads
$\mathrm{d}Z_{t}=-A_{\text{{cPD}}}Z_{t}\mathrm{d}t+B_{\text{{cPD}}}\mathrm{d}W_{t},\quad\text{with }A_{\text{{cPD}}}=\begin{pmatrix}c_{g}^{-1}&k\\ -\lambda k&\lambda c_{f}\end{pmatrix},B_{\text{{cPD}}}=\begin{pmatrix}\sqrt{2}\\ 0\end{pmatrix}$
where $W_{t}$ is one-dimensional Brownian motion and the weight $\lambda>0$ relates primal and dual time scales. Denoting $E(t):=(e_{ij}(t))_{i,j}:=\exp(-tA_{\text{{cPD}}}^{T})$, the following statements hold:
@@
    If $\pi_{0}$ is Gaussian (including degenerate Gaussians with singular covariance matrix), then $\pi_{t}$ is (non-degenerate) Gaussian for all $t>0$. As $t\to\infty$, the solution $\pi_{t}$ converges to the Gaussian stationary $\pi^{\text{{stat}}}$ with mean zero and covariance matrix
@@
    The $x$-marginal $\mu^{\text{{stat}}}_{\text{{cPD}}}$ is not equal to the target $\mu^{\ast}$. However, the limiting joint distribution as $\lambda\to\infty$ is the degenerate Gaussian $(id,f^{\prime}(k\,\cdot))_{\#}\mu^{\ast}$ whose $x$-marginal is the target $\mu^{\ast}$.
@@
$\partial_{t}\pi_{t}=\nabla_{z}\cdot(\pi_{t}A_{\text{{cPD}}}z)+\Delta_{x}\pi_{t},$
@@
$\left\{\begin{array}[]{l}\partial_{t}p_{t}=-\zeta^{T}A_{\text{{cPD}}}\nabla_{\zeta}p_{t}-\xi^{2}p_{t}\\ p_{0}(\xi,\omega)=\mathcal{F}\pi_{0}.\end{array}\right.$
@@
$\dot{\zeta}(t)$ $=A_{\text{{cPD}}}^{T}\zeta(t)$
@@
$\zeta(t)$ $=\exp(tA_{\text{{cPD}}}^{T})\zeta_{0}$
@@
By eliminating the initial value $\zeta_{0}=\exp(-tA_{\text{{cPD}}}^{T})\zeta(t)$, the solution for the initial value problem is
$p_{t}(\zeta)=p_{0}(\exp(-tA_{\text{{cPD}}}^{T})\zeta)\exp\left(-\frac{1}{2}\int_{0}^{t}(B_{\text{{cPD}}}^{T}\exp(-sA_{\text{{cPD}}}^{T})\zeta)^{2}\,\mathrm{d}s\right).$
The quadratic form inside the integral is positive semidefinite in the variable $\zeta$. For $t>0$, it is also positive definite: Otherwise, there would need to be a $\zeta$ such that $B_{\text{{cPD}}}^{T}\exp(-sA_{\text{{cPD}}}^{T})\zeta=0$ for all $s\in[0,t]$, which implies $B_{\text{{cPD}}}^{T}(A_{\text{{cPD}}}^{T})^{k}\zeta=0$ for all $k\geq 0$. This is equivalent to $(0,1)^{T}$ being an element of an invariant subspace of $A_{\text{{cPD}}}^{T}$, which is impossible since the eigenvectors of $A_{\text{{cPD}}}^{T}$ are of the form
@@
If the initialization $\pi_{0}$ is Gaussian with mean $z_{0}=(x_{0},y_{0})$ and (possibly singular) covariance $\Sigma_{0}$, we can simply take the inverse Fourier transform to obtain the solution $\pi_{t}$. It is worth noting, however, that we do not even have to compute the full distribution $\pi_{t}$ in order to analyze the $x$-marginal $\mu_{t}$ of $\pi_{t}$. Since marginalizing out the dual variable $y$ in $\pi_{t}$ corresponds to evaluating $p_{t}$ at $\zeta=(\xi,0)$ in Fourier space, we get, denoting by $\mathcal{F}^{-1}_{\xi}$ the inverse Fourier transform in $\xi$ and $E(t):=(e_{ij}(t))_{i,j}:=\exp(-tA_{\text{{cPD}}}^{T})$,

### v3

@@ 2.2 Continuous Time Perspective
In order to analyze ULPDA, we want to define a similar time-continuous limit of (1.4). We assume that the step sizes $\tau,\sigma$ are coupled via the constant $\lim_{\tau,\sigma\to 0}\frac{\sigma}{\tau}=\lambda$ where $0<\lambda<\infty$. Taking both step sizes to zero at this ratio, the formal limiting Markov process $(X_{t},Y_{t})$ satisfies the diffusion equation
$\left\{\begin{array}[]{l}\,\mathrm{d}X_{t}\in-(\partial g(X_{t})+K^{T}Y_{t})\,\mathrm{d}t+\sqrt{2}\,\mathrm{d}W_{t}\\ \,\mathrm{d}Y_{t}\in-\lambda(\partial f^{*}(Y_{t})-KX_{t})\,\mathrm{d}t,\end{array}\right.$ (2.5)
@@
$\,\mathrm{d}Z_{t}\in-A_{\text{{PD}}}(Z_{t})\,\mathrm{d}t+B_{\text{{PD}}}\,\mathrm{d}W_{t}.$ (2.6)
where we introduced the primal-dual coefficients
$A_{\text{{PD}}}(Z_{t})\coloneqq\begin{pmatrix}\partial g(X_{t})+K^{T}Y_{t}\\ -\lambda(KX_{t}+\partial f^{\ast}(Y_{t}))\end{pmatrix},\qquad B_{\text{{PD}}}\coloneqq\begin{pmatrix}\sqrt{2}I_{d}\\ 0\end{pmatrix}.$ (2.7)
Note in particular that the diffusion coefficient $\Sigma_{\text{{PD}}}\coloneqq\frac{1}{2}B_{\text{{PD}}}B_{\text{{PD}}}^{T}$ is only positive semi-definite (or degenerate) with an $m$-dimensional null space, where $m$ is the dimension of the dual variable $Y$. Standard existence and uniqueness analysis for kinetic Fokker-Planck equations with positive definite diffusion coefficient does not hold anymore.
In the more general scheme (2.4), the diffusion equation takes the form
$\,\mathrm{d}Z_{t}\in-A_{\text{{PD}}}(Z_{t})\,\mathrm{d}t+B_{\text{{gPD}}}\,\mathrm{d}W_{t},\qquad B_{\text{{gPD}}}\coloneqq\begin{pmatrix}B_{X}\\ B_{Y}\end{pmatrix}.$ (2.8)
If $B_{\text{{gPD}}}$ (where ‘gPD’ denotes ‘general primal-dual’) is invertible, then $\Sigma_{\text{{gPD}}}\coloneqq\frac{1}{2}B_{\text{{gPD}}}B_{\text{{gPD}}}^{T}$ becomes positive definite, and the corresponding Fokker-Planck equation has a solution [Pavliotis2014]. We will check whether the stationary solution can be related to the density of the target distribution $\mu^{\ast}$.
3 The Stationary Solution in Continuous Time
@@
$\mathcal{T}(\mu,\tilde{\mu}):=\inf_{\mathcal{M}\in\Gamma(\mu,\tilde{\mu})}\left(\mathbb{E}_{(x,\tilde{x})\sim\mathcal{M}}\left[\lVert x-\tilde{x}\rVert^{2}\right]\right)^{1/2},$ (3.1)
where $\Gamma(\mu,\tilde{\mu})$ denotes the set of all couplings with marginals $\mu$ and $\tilde{\mu}$, i.e.
@@ 3.1 Fokker-Planck Equation and Existence of a Stationary Solution
The contraction result follows an argument from the literature on overdamped Langevin diffusion, see [Fournier2021, Durmus2019]. Since this simpler case is instructive, we repeat it here and then generalize to the setting of an SDE with any strongly monotone operator as drift coefficient.
3.1.1 Contraction of Overdamped Langevin Diffusion via a Coupling Argument
Suppose for now that the potential $h:\mathbb{R}^{d}\to\mathbb{R}$ is differentiable and $\omega_{h}$-strongly convex. We repeat here an instructive calculation that has been carried out, e.g., in [Fournier2021, Durmus2019] to show stability of the overdamped Langevin diffusion process
@@ 3.1.1 Contraction of Overdamped Langevin Diffusion via a Coupling Argument
In particular, if we choose one of the processes to be initialized at the target distribution as $\tilde{\mu}_{0}=\mu_{\ast}$, then the distribution of $\tilde{X}_{t}$ is stationary at $\tilde{\mu}_{t}=\mu^{\ast}$ and by the above we obtain the convergence result
@@ 3.1.2 Contraction and Stationary Solution of the Primal-Dual Diffusion Equation
We can generalize the above argument allowing an analysis of the time-continuous limit (2.6) of the primal-dual sampling scheme (1.4). Although the primal-dual dynamics are not the gradient flow of a corresponding relative entropy anymore, contraction can equivalently be shown for a coupling of two processes. For that, let $Z_{t}$ satisfy a general stochastic differential inclusion with non-homogeneous coefficients
@@
Here $W_{t}$ is Brownian motion in $\mathbb{R}^{K}$ and $B:\mathbb{R}^{d}\to\mathbb{R}^{d\times K}$. We assume the drift $A$ to be a set-valued monotone operator, meaning
@@
with $\Sigma(z)=\frac{1}{2}B(z)B(z)^{T}$.
@@
$\begin{pmatrix}\mathrm{d}Z_{t}\\ \mathrm{d}\tilde{Z}_{t}\end{pmatrix}\in-\begin{pmatrix}A(Z_{t})\\ A(\tilde{Z}_{t})\end{pmatrix}\mathrm{d}t+\begin{pmatrix}B(Z_{t})\\ B(\tilde{Z}_{t})\end{pmatrix}\mathrm{d}W_{t},$ (3.6)
@@
$\frac{\partial\Pi_{t}}{\partial t}=\nabla_{z}\cdot\left(w\pi_{t}\right)+\nabla_{\tilde{z}}\cdot\left(\tilde{w}\Pi_{t}\right)+(\nabla_{z}+\nabla_{\tilde{z}})\cdot(\Sigma(z,\tilde{z})(\nabla_{z}+\nabla_{\tilde{z}})\Pi_{t}),\qquad w\in A(z),\tilde{w}\in A(\tilde{z}).$ (3.7)
@@
With a change of variable we can write $\nabla_{z}+\nabla_{\tilde{z}}=\nabla_{z+\tilde{z}}$. By employing (3.7), we obtain the same type of inequality as in the case of overdamped Langevin diffusion that
@@
$=\frac{1}{2}\int\lVert z-\tilde{z}\rVert^{2}\left(\nabla_{z}\cdot\left(w\Pi_{t}\right)+\nabla_{\tilde{z}}\cdot\left(\tilde{w}\Pi_{t}\right)+(\nabla_{z+\tilde{z}})\cdot(\Sigma(\nabla_{z+\tilde{z}})\Pi_{t})\right)\,\mathrm{d}z\,\mathrm{d}\tilde{z}$
@@
Consider now the general primal-dual type diffusion (2.8) and denote $\pi_{t}$ for the law of $Z_{t}$. By Proposition 3.1, $\pi_{t}$ solves the Fokker-Planck equation
$\frac{\partial\pi_{t}}{\partial t}=\nabla_{z}\cdot\left(w\pi_{t}+\Sigma_{\text{{gPD}}}\nabla_{z}\pi_{t}\right),\qquad w\in A_{\text{{PD}}}(z).$ (3.8)
with $\Sigma_{\text{{gPD}}}=\frac{1}{2}B_{\text{{gPD}}}B_{\text{{gPD}}}^{T}$ as in (2.8) and $A_{\text{{PD}}}$ given in (2.7). The drift coefficient $A_{\text{{PD}}}$ is a monotone operator.
[…]
The transport distance (3.1) on $\mathcal{P}_{2}(\mathcal{Z})$ induced by $\lVert\cdot\rVert_{a,b}$ will be denoted by $\mathcal{T}_{a,b}$. Note that the equivalence of norms on $\mathbb{R}^{n}$ implies equivalence of all $\mathcal{T}_{a,b}$ for $a,b>0$.
@@
Assume that $g,f^{\ast}$ are $\omega_{g}$-strongly convex and $\omega_{f^{\ast}}$-strongly convex, respectively, with $\omega_{g},\omega_{f^{\ast}}>0$. Let $Z_{t},\tilde{Z}_{t}$ be two coupled processes that solve the general primal-dual diffusion equation (2.8) with corresponding densities $\pi_{t}$ and $\tilde{\pi}_{t}$. Then it holds
@@
where $\omega=\min\{\omega_{g},\lambda\omega_{f^{\ast}}\}>0$. In particular, the dynamics are contractive and there exists a unique stationary distribution of the continuous time primal-dual diffusion. Making the dependence on $\lambda$ explicit, we denote this stationary distribution by $\pi_{\lambda}$.
@@
We employ Lemma 3.3, hence the only part we need to show is the strong monotonicity of the negative drift coefficient $A_{\text{{PD}}}$ defined by
$A_{\text{{PD}}}=\begin{pmatrix}\partial g&K^{T}\\ -\lambda K&\lambda\partial f^{\ast}\end{pmatrix}.$
We write $z=(x,y)\in\mathcal{X}\times\mathcal{X}=\mathcal{Z}$ for the joint primal and dual variable and equip $\mathcal{Z}$ with the weighted inner product $\langle\cdot,\cdot\rangle_{1,\lambda^{-1}}$. Let now $w\in A_{\text{{PD}}}(z)$ and $\tilde{w}\in A_{\text{{PD}}}(\tilde{z})$. Then there are $u\in\partial g(x)$ and $v\in\partial f^{\ast}(y)$ such that $w=(u+K^{T}y,\lambda(v-Kx))$. Equivalently, if $\tilde{w}\in A(\tilde{z})$ then there exist $\tilde{u}\in\partial g(\tilde{x})$ and $\tilde{v}\in\partial f^{\ast}(\tilde{y})$ with $\tilde{w}=(\tilde{u}+K^{T}\tilde{y},\lambda(\tilde{v}-K\tilde{x}))$. It follows that
@@
The stationary solution of (3.8) will be denoted by $\pi_{\lambda}$. We will also use the $x$-marginal $\mu_{\lambda}$ of $\pi_{\lambda}$, defined as $\mu_{\lambda}(x)\coloneqq\int_{\mathcal{Y}}\pi_{\lambda}(x,y)\,\mathrm{d}y.$
@@ 3.2 Non-Concentration of the Stationary Solution
Having established the existence of $\pi_{\lambda}$, we can try to relate the algorithmic, time discretized scheme to it. Start by considering the special case (2.6) where diffusion is only added in the primal variable. While overdamped Langevin dynamics is designed to have the target $\mu^{\ast}$ as its stationary solution, this is not the case in the primal-dual setting. In general, $\mu^{\ast}\neq\mu_{\lambda}$, in other words $\mu^{\ast}$ is not the $x$-marginal of $\pi_{\lambda}$. We demonstrate this on a simple example of a Gaussian target, where we can solve the Fokker-Planck equation (3.8) in closed form.
@@
where $k\neq 0$ and $c_{g},c_{f}>0$. It holds $f^{\ast}(y)=\frac{c_{f}}{2}y^{2}$ for all $y\in\mathbb{R}$. The primal-dual diffusion equation (2.6) for the joint variable $Z_{t}=(X_{t},Y_{t})^{T}\in\mathbb{R}^{2}$ reads
$\mathrm{d}Z_{t}=-A_{\text{{PD}}}Z_{t}\mathrm{d}t+B_{\text{{PD}}}\mathrm{d}W_{t},\quad\text{with }A_{\text{{PD}}}=\begin{pmatrix}c_{g}^{-1}&k\\ -\lambda k&\lambda c_{f}\end{pmatrix},B_{\text{{PD}}}=\begin{pmatrix}\sqrt{2}\\ 0\end{pmatrix}$
where $W_{t}$ is one-dimensional Brownian motion and the weight $\lambda>0$ relates primal and dual time scales. Denoting $E(t):=(e_{ij}(t))_{i,j}:=\exp(-tA_{\text{{PD}}}^{T})$, the following statements hold:
@@
    If $\pi_{0}$ is Gaussian (including degenerate Gaussians with singular covariance matrix), then $\pi_{t}$ is (non-degenerate) Gaussian for all $t>0$. As $t\to\infty$, the solution $\pi_{t}$ converges to the Gaussian stationary $\pi_{\lambda}$ with mean zero and covariance matrix
@@
    The $x$-marginal $\mu_{\lambda}$ is not equal to the target $\mu^{\ast}$. However, the limiting joint distribution as $\lambda\to\infty$ is the degenerate Gaussian $(id,f^{\prime}(k\,\cdot))_{\#}\mu^{\ast}$ whose $x$-marginal is the target $\mu^{\ast}$.
@@
$\partial_{t}\pi_{t}=\nabla_{z}\cdot(\pi_{t}A_{\text{{PD}}}z)+\Delta_{x}\pi_{t},$
@@
$\left\{\begin{array}[]{l}\partial_{t}p_{t}=-\zeta^{T}A_{\text{{PD}}}\nabla_{\zeta}p_{t}-\xi^{2}p_{t}\\ p_{0}(\xi,\omega)=\mathcal{F}\pi_{0}.\end{array}\right.$
@@
$\dot{\zeta}(t)$ $=A_{\text{{PD}}}^{T}\zeta(t)$
@@
$\zeta(t)$ $=\exp(tA_{\text{{PD}}}^{T})\zeta_{0}$
@@
By eliminating the initial value $\zeta_{0}=\exp(-tA_{\text{{PD}}}^{T})\zeta(t)$, the solution for the initial value problem is
$p_{t}(\zeta)=p_{0}(\exp(-tA_{\text{{PD}}}^{T})\zeta)\exp\left(-\frac{1}{2}\int_{0}^{t}(B_{\text{{PD}}}^{T}\exp(-sA_{\text{{PD}}}^{T})\zeta)^{2}\,\mathrm{d}s\right).$
The quadratic form inside the integral is positive semidefinite in the variable $\zeta$. For $t>0$, it is also positive definite: Otherwise, there would need to be a $\zeta$ such that $B_{\text{{PD}}}^{T}\exp(-sA_{\text{{PD}}}^{T})\zeta=0$ for all $s\in[0,t]$, which implies $B_{\text{{PD}}}^{T}(A_{\text{{PD}}}^{T})^{k}\zeta=0$ for all $k\geq 0$. This is equivalent to $(0,1)^{T}$ being an element of an invariant subspace of $A_{\text{{PD}}}^{T}$, which is impossible since the eigenvectors of $A_{\text{{PD}}}^{T}$ are of the form
@@
If the initialization $\pi_{0}$ is Gaussian with mean $z_{0}=(x_{0},y_{0})$ and (possibly singular) covariance $\Sigma_{0}$, we can simply take the inverse Fourier transform to obtain the solution $\pi_{t}$. It is worth noting, however, that we do not even have to compute the full distribution $\pi_{t}$ in order to analyze the $x$-marginal $\mu_{t}$ of $\pi_{t}$. Since marginalizing out the dual variable $y$ in $\pi_{t}$ corresponds to evaluating $p_{t}$ at $\zeta=(\xi,0)$ in Fourier space, we get, denoting by $\mathcal{F}^{-1}_{\xi}$ the inverse Fourier transform in $\xi$ and $E(t):=(e_{ij}(t))_{i,j}:=\exp(-tA_{\text{{PD}}}^{T})$,